Using mathematical induction, prove that for every positive integer n,

1 · 2 + 2 · 3 + · · · + n(n + 1) = (n(n + 1)(n + 2)) / 3 


1
Expert's answer
2021-10-26T15:21:11-0400

Base step

for n=1: "1*2=\frac{1*2*3}{3}=2." The statement holds.

Inductive step

Assume that for n=k: "1*2+2*3+...+k(k+1)=\frac{k(k+1)(k+2)}{3}"

Then, for n=k+1: "1*2+2*3+...+k(k+1)+(k+1)(k+2)=\frac{k(k+1)(k+2)}{3}+(k+1)(k+2)="

"=(k+1)(k+2)(\frac{k}{3}+1)=\frac{(k+1)(k+2)(k+3)}{3}"

That is, the statement for n=k+1 also holds true.

Since both the base step and the inductive step have been proved as true, by mathematical induction the statement holds for every natural number n.